OpenStudy (anonymous):

Evaluate the function T(x) =19 -x, when x =-3

OpenStudy (anonymous):

For all functions involving x, just substitute x into the function. So if \(T(x)=19-x\) when \(x=-3\), the function simply becomes \(T(-3)=19-(-3)\)

OpenStudy (anonymous):

T(x) =19 -x put x = -3 here and get your answer